# Taxonomy of periodic orbits and gravitational waves in a deformed Schwarzschild black hole spacetime

### v2

@@ Abstract
In this paper, we investigate periodic orbits of test particles around a deformed Schwarzschild black hole and the resulting gravitational waves. Firstly, we examine the properties of circular orbits and find that circular orbits could disappear when the deformation is large enough. Then, using an orbital taxonomy, we characterize various periodic orbits with a set of triples, which describes the zoom-whirl behaviours. We also calculate the gravitational waveform signals generated by different periodic orbits, revealing the influence of the deformation on the gravitational wave, which can be potentially picked up by future space-based detectors.
@@ 1 Introduction
Since the successful detection of gravitational wave signal GW150914 in 2016 by LIGO and Virgo, gravitational waves have been a new messenger of cosmic explorationAbbott and others (2016b, c, a, d). These oscillations of spacetime were predicted by Einstein’s theory of general relativity and have significantly broadened the research horizon on the physics frontierBarack and others (2019); Berti and others (2015). Among all the sources for space-based gravitational wave detectors——such as TaijiHu and Wu (2017), TianqinLuo and others (2016); Gong et al. (2021), and LISASchutz (1999); Amaro-Seoane et al. (2017); Gair et al. (2004)——the Extreme Mass Ratio Inspiral system (EMRI) is a crucial class of gravitational wave sources in the mHz frequency band, in which a stellar-mass object inspirals around a supermassive black holeHughes (2001).
@@ 5 Gravitational Radiation from periodic orbits
In order to show the influence of the deformation parameter on the gravitational waveforms of different periodic orbits, we here consider a small compact object with mass $m=10M_{\odot}$ around a supermassive black hole with mass $M=10^{6}M_{\odot}$ at the luminosity distance $D_{\text{L}}=1\text{Gpc}$. Take a simple case, the inclination angle $\iota$ is set as $\pi/2$ and the longitude of perihelion is fixed at 0. Moreover, the polarization angle is also set as 0 for simplicityBabak et al. (2007). To show contrast, we fix the energy at $E=\sqrt{0.9755751102284231}$ and change the deviation parameter $\alpha$. When $\alpha$ is large enough, the perihelion is too small, as shown in Fig.6. So numerical Kludge model will not work. And then the waveforms are represented by two polarization components $h_{+}$ and $h_{\times}$ to show their characteristics in Fig.7. We list three parameter conditions and their waveform contrast here.
@@
Figure 7: Two polarized components of gravitational waveforms of periodic orbit from a test object with $m=10M_{\odot}$ around a supermassive black hole with $M=10^{6}M_{\odot}$. Energy is fixed as $E=\sqrt{0.9755751102284231}$. The above figure shows the difference between $\alpha=0$ and $\alpha=0.2$. The below figure shows the difference between $\alpha=0$ and $\alpha=0.4$.
In Figs6 and 7, a typical kind of periodic orbits is plotted with the triple $(1,2,0)$ and their $h_{+}$ and $h_{\times}$ modes with different deformation parameters. It shows that a clear relationship exists between gravitational waveforms and orbit. The giant leave results in very long smooth phases. And there are rapid oscillations which align with the zoom and whirl behaviors. Moreover, it is clear that in comparison with the Schwarzschild case the parameter influences the phase of the waveforms mainly and the waveform amplitudes in a slight way when parameter increases. To clearly illustrate the differences in gravitational waveforms with different parameters, a useful characteristic is the mismatch between to waveformsAllen (2019); Qiao et al. (2025); Lai et al. (2025). To conduct a preliminary analysis of the differences, we calculated the mismatch according to the Schwarzschild condition with $\alpha$ from 0 to 0.4 and concluded the trend below in Fig.8:
Figure 8: Mismatch based on the Schwarzschild case with the parameter $\alpha$ from 0 to 0.4.
We find that the mismatch will be over $0.1$ when $\alpha\geq 0.5$, so we don’t consider it. The value of mismatch, apparently, goes up with the increase of the deformation parameter. It is reasonable because it means that difference from the Schwarzschild situation increases. We can preliminarily distinguish the difference by the value of mismatch accumulating over time although the effect of the deformation is relatively weak in the amplitude and phases. Certainly, further detailed differentiation of gravitational waveforms needs more detailed detectability and measurability analysis. Further research can be conducted based on the reference to Speeney et al. (2022). These findings have pointed the way of detecting detailed physical characteristics in the future.

### v3

@@ Abstract
In this paper, we investigate periodic orbits of test particles around a deformed Schwarzschild black hole and the resulting gravitational waves. Firstly, we examine the properties of circular orbits and find that circular orbits could disappear when the deformation is large enough. Then, using an orbital taxonomy, we characterize various periodic orbits with a triplet of integers, which describes the zoom-whirl behaviours. We also calculate the gravitational waveform signals generated by different periodic orbits, revealing the influence of the deformation on the gravitational wave, which can be potentially picked up by future space-based detectors.
@@ 1 Introduction
Since the successful detection of the gravitational-wave signal GW150914 in 2016 by LIGO and Virgo, gravitational waves have become a new messenger of cosmic explorationAbbott and others (2016b, c, a, d). These oscillations of spacetime were predicted by Einstein’s theory of general relativity and have significantly broadened the research horizon on the physics frontierBarack and others (2019); Berti and others (2015). Among all the sources for space-based gravitational wave detectors——such as TaijiHu and Wu (2017), TianqinLuo and others (2016); Gong et al. (2021), and LISASchutz (1999); Amaro-Seoane et al. (2017); Gair et al. (2004)——the Extreme Mass Ratio Inspiral system (EMRI) is a crucial class of gravitational wave sources in the mHz frequency band, in which a stellar-mass object inspirals around a supermassive black holeHughes (2001).
@@ 5 Gravitational Radiation from periodic orbits
In order to show the influence of the deformation parameter on the gravitational waveforms of different periodic orbits, we consider a small compact object with mass $m=10M_{\odot}$ around a supermassive black hole with mass $M=10^{6}M_{\odot}$ at the luminosity distance $D_{\text{L}}=1\text{Gpc}$. Take a simple case, the inclination angle $\iota$ is set as $\pi/2$ and the longitude of perihelion is fixed at 0. Moreover, the polarization angle is also set as 0 for simplicityBabak et al. (2007). To show contrast, we fix the energy at $E=\sqrt{0.9755751102284231}$ and change the deviation parameter $\alpha$. When $\alpha$ is large enough, the perihelion is too small, as shown in Fig.6. So numerical kludge model will not work. And then the waveforms are represented by two polarization components $h_{+}$ and $h_{\times}$ to show their characteristics in Fig.7. We list three parameter conditions and their waveform contrast here.
@@
Figure 7: Two polarization components of gravitational waveforms of periodic orbit from a test object with $m=10M_{\odot}$ around a supermassive black hole with $M=10^{6}M_{\odot}$. Energy is fixed as $E=\sqrt{0.9755751102284231}$. The above figure shows the difference between $\alpha=0$ and $\alpha=0.2$. The below figure shows the difference between $\alpha=0$ and $\alpha=0.4$.
In Figs 6 and 7, a typical kind of periodic orbits is plotted with the triplet $(1,2,0)$ and their $h_{+}$ and $h_{\times}$ modes with different deformation parameters. It shows that a clear relationship exists between gravitational waveforms and orbit. The giant leaf results in very long smooth phases. And there are rapid oscillations which align with the zoom and whirl behaviors. Moreover, it is clear that in comparison with the Schwarzschild case the parameter influences the phase of the waveforms mainly and the waveform amplitudes in a slight way when parameter increases. To clearly quantitatively illustrate the differences in gravitational waveforms with different parameters, a useful characteristic is the mismatch between two waveformsAllen (2019); Qiao et al. (2025); Lai et al. (2025). The specific calculation formula is given asPompili (2025):
$\mathcal{M}=1-\max_{\varphi,t_{c}}\frac{(h_{1}\mid h_{2})}{\sqrt{(h_{1}\mid h_{1})(h_{2}\mid h_{2})}}$ (29)
While $h$ is defined as $h=h_{+}-ih_{\times}$. To conduct a preliminary analysis of the differences, we calculated the mismatch according to the Schwarzschild condition with $\alpha$ from 0 to 0.38 and concluded the trend below in Fig.8. To show this trend more comprehensively, we display mismatch of three different periodic orbits, with the triplet $(1,2,0)$, $(2,1,1)$ and $(3,1,1)$, respectively. We display the smooth curve with the marked original data points in the figure.:
Figure 8: Mismatch based on the Schwarzschild case with the parameter $\alpha$ from 0 to 0.38. The blue, red and yellow line shows the mismatch of orbits with the triplet $(1,2,0)$, $(2,1,1)$ and $(3,1,1)$, respectively. The bolded points represent the original data.
Although not shown in Fig. 8 , we find that the mismatch exceeds $0.05$ for $\alpha\geq 0.4$. The mismatch increases with the deformation parameter, as expected for progressively larger departures from the Schwarzschild situation. We also find that the mismatch varies across different periodic orbits, indicating an orbit-dependent sensitivity to the deformation. It is noteworthy that with the same number of $w$ and $v$, the mismatch of the orbit with three branches is greater than that of the orbit with two branches, which may be attributed to the accumulation of waveform differences induced by the number of orbital branches. While the deformation induces only modest changes in the waveform amplitude and phase, the mismatch provides a preliminary means of discriminating among these cases. A definitive assessment, however, requires a dedicated analysis of detectability and measurability, following and extending the methodology of Ref.Speeney et al. (2022). Overall, these results suggest a plausible pathway toward probing subtle physical features in future gravitational wave detection.